# Solid-State Dewetting of Polycrystalline Thin Films: a Phase Field Approach

### v1

@@ Abstract
Solid-state dewetting (SSD) is the process by which thin solid films break up and retract on a substrate, resulting in the formation of nanostructures and islands. While SSD in single-crystalline films is generally understood as a surface-energy-driven process mediated by surface diffusion, polycrystalline films feature additional complexity due to the presence of grain boundaries. Theoretical investigations and simulation frameworks have mainly focused on single-crystalline SSD. In this work, we present and apply a multi-phase-field framework that captures key mechanisms in polycrystalline thin films with multiple grains and grain boundaries in three dimensions. By considering isotropic surface and interface energy and curvature-driven grain boundary motion, we demonstrate how the framework reproduces the key phenomenology of SSD while being consistent with predictions based on energetic arguments. We also introduce refined analytical criteria for the onset of dewetting, confirmed by numerical simulations. Key evidence is given concerning the onset of dewetting at grain-boundary triple junctions. Moreover, we showcase SSD of polycrystalline patches. This work paves the way for in-depth investigations of the morphological evolution of polycrystalline thin films.
@@
Thin crystalline films are metastable due to their high surface-to-volume ratio. At high temperatures below melting, they typically break up and form three-dimensional agglomerates in a process known as solid-state dewetting (SSD) [1]. While detrimental to planar architectures, this phenomenon realizes a unique kinetic pathway for the self-assembly of complex nanostructures. Various strategies have been proposed to control the outcomes of SSD in single-crystalline thin films, such as patterning the film or the underlying substrate [2, 3, 4, 5]. In these systems, the dynamics are predominantly governed by surface diffusion [1].
@@
Several technologically relevant thin films are polycrystalline [6, 7]: they consist of crystalline domains (grains) with varying crystallographic orientations, which are separated by interfaces known as grain boundaries (GBs). Their morphological evolution is influenced not only by surface diffusion, as in single crystals, but also by additional effects associated with GBs [8, 9, 10, 11]. Arguably, the most relevant one is that surface grooves develop at GBs. Moreover, GBs migrate through a mechanism resembling attachment-detachment kinetics, with a velocity governed largely by capillarity (mean curvature flow) and potentially influenced by elasticity effects [12, 13].
@@
Modeling and numerical simulation played a critical role in understanding and tailoring single-crystalline SSD (see, e.g., [14, 4, 5]). Studies on GB grooving and few-grains microstructures have also yielded valuable insights [15, 16, 17, 18, 19], but remain limited in capturing the coupled dynamics of GB migration and surface diffusion across many grains and long timescales. Phase field (PF) models [20, 21, 22, 23] hold a prominent position to study SSD. They realize an implicit description of geometries via smooth order parameter(s), which naturally deals with topological changes, complex geometries, and three-dimensionality, all key aspects of SSD [24, 4, 5, 25, 26]. Convenient PF models coupling diffusion, phase transformation, and interface migration have been proposed [27, 28, 18, 29, 30]. A recent extension of the grand-potential PF model [27, 28, 31] accurately captures bulk and surface diffusion coupled with interface migration, reproducing key features of polycrystalline SSD, such as thermal grooving at GBs [29].
In this work, we present a multi-PF (MPF) approach to studying SSD of polycrystalline thin films. After a brief overview of the model’s key features, we present its capabilities and validate it through benchmarks. Furthermore, we discuss novel insights from simulations and semi-analytic equations regarding the onset of SSD. We also apply the MPF model to prototype polycrystalline films and patches, outlining differences and similarities with their single-crystalline counterparts.
@@
We employ the MPF model introduced in [29], considering its specialization to systems dominated by surface/interface diffusion, as well as interface migration, while neglecting bulk diffusion. We also consider grains of the same pure material and neglect any elasticity effects that may arise from GB migration. In brief, order parameters $\{\phi_{1}(\mathbf{r}),\dots,\phi_{N}(\mathbf{r})\}\equiv\boldsymbol{\varphi}$ akin to smooth characteristic functions track different domains/phases. To account for material transport, global concentration fields for $K$ elements $c_{i}(\mathbf{r})$ are considered. They can be related to the composition of each phase ($c_{i\alpha}$) via interpolation $c_{i}(\mathbf{r})=\sum_{\alpha=1}^{N}\phi_{\alpha}(\mathbf{r})c_{i\alpha}$ (spatial dependence will be omitted for brevity). In order to describe the chemical interactions within the MPF system, we define a set of $K-1$ independent chemical potential fields, $\{\mu_{1},\dots,\mu_{K-1}\}\equiv\boldsymbol{\mu}$, corresponding to the concentration fields, which distinguish between the solid film, the substrate, and the surrounding vacuum. The grand potential functional reads
$\begin{split}\Psi[\boldsymbol{\varphi},\nabla\boldsymbol{\varphi},\boldsymbol{\mu}]=&\int_{\Omega}\bigg{[}f_{\rm I}(\boldsymbol{\varphi},\nabla\boldsymbol{\varphi})+\psi(\boldsymbol{\varphi},\boldsymbol{\mu})\bigg{]}\mathrm{d}\mathbf{r},\\ \end{split}$ (1)
@@
$f_{\rm I}=\sum_{\alpha}^{N}\sum_{\beta=\alpha+1}^{N}\gamma_{\alpha\beta}\left(\epsilon\nabla\phi_{\alpha}\cdot\nabla\phi_{\beta}+\frac{16\mathcal{G}(\boldsymbol{\varphi})\phi_{\alpha}\phi_{\beta}}{\pi^{2}\epsilon}\right),$ (2)
encodes the interfacial energy with $\epsilon$ the interface thickness, $\mathcal{G}(\boldsymbol{\varphi})=1$ if $\sum_{\alpha}\phi_{\alpha}=1$ and $\phi_{\alpha}\geq 0,\ \forall\alpha$, while $\mathcal{G}(\boldsymbol{\varphi})=\infty$ otherwise (realizing a double-obstacle potential). The function $\psi(\boldsymbol{\varphi},\boldsymbol{\mu})$ corresponds to the grand-potential density
@@
In this study we used a parabolic free energy density $f_{i\alpha}=A_{i}^{\rm chem}(c_{i\alpha}-c_{i\alpha}^{\text{eq}})^{2}$ depending on the equilibrium concentrations $c_{i\alpha}^{\text{eq}}$ and constants $A_{i}^{\rm chem}$.
@@
with $m_{\alpha\beta}$ and $M_{\alpha\beta}$ the mobility and diffusivity at the $\alpha\beta$-interface, respectively. The timescale will be expressed in arbitrary units. Sharp interface limits showing convergence to the targeted kinetics (surface diffusion coupled with attachment-detachment kinetics) can be found in [29, 32]. Simulations are obtained by numerically solving Eqs. (4) using a finite difference integration scheme, implemented in the modular software package for multi-physics, parallel computations PACE3D [33]. Periodic boundary conditions are applied in the in-plane directions, while no-flux Neumann conditions are imposed at the top and bottom boundaries.
@@
Figure 1: Modeling and 2D simulation of polycrystalline SSD. (a) Scheme of the considered geometries and labeling of the phases as considered in the MPF model. The grain aspect ratio is defined as $r^{\rm 2D}=W/H$. (b) SSD of a thin film with three grains. The central one has an aspect ratio of $r^{\rm 2D}=25$. The two at the border of the film have both $r^{\rm 2D}=12.5$. The timescale is relative to the one needed to reach equilibrium. (c) MPF simulations of SSD with varying $r^{\rm 2D}$ for a film ideally formed by grains of the same size. A critical aspect ratio of $r^{\rm 2D}_{\rm c}\approx 9$ is obtained in simulations.
@@
Figure 2: Geometrical parametrization and estimation of the critical aspect ratios. (a) Schematics of 2D equilibrium morphology. The dashed line illustrates the thickness ($H$) of the conformal initial film. (b) Schematics (top view) of the idealized setting of hexagonal grains in 3D with equal size. (c) Spherical caps truncated by $n$-gons. (d) $r_{\rm c}(\theta)$ in 2D and 3D as predicted by Eq. (5) and Eq. (8) for $d=0$ (sharp interface limit) and $d>0$ as realized in the MPF simulations. The vertical dotted line indicates the $\theta$ realized in the MPF simulations. Symbols show $r_{\rm c}$ obtained in 2D (circle) and 3D (square) MPF simulations.
@@
In our polycrystalline thin film model, the order parameters $\phi_{\alpha}$ track the substrate, the vapor phases, and the different grains within the film. Fig. 1a shows an illustration of a polycrystalline film composed of three grains, where we label the vacuum phase $\alpha=1$, the substrate $\alpha=N$, and the grains $1<\alpha<N$. We refer to the ratio of width to height $r^{\rm 2D}=W/H$ as the aspect ratio of the 2D grains. We use a two-concentration-field model for the substrate and the film with $c_{\text{substrate},N}^{\text{eq}}=1$ and $c_{\text{grain},\alpha}^{\text{eq}}=1$. We set $A_{\text{substrate}}^{\rm chem}=50$ and $A_{\text{grain}}^{\rm chem}=5$ [32]. Interfacial (surface) diffusion is only allowed at the film-vapor interface via $M_{\alpha\beta}=\delta_{1\beta}$. GB mobilities are set to $10^{-3}$ while all other mobilities set to $1$ and $\gamma_{\alpha\beta}=0.5$ for all interfaces.
@@
A first representative simulation of SSD for a 2D three-grains film, with aspect ratios $r^{\rm 2D}_{\text{left}}=r_{\text{right}}^{\rm 2D}=12.5$ and $r_{\text{center}}^{\rm 2D}=25$, is shown in Fig. 1b. This setting approximates stripes infinitely extended in the direction perpendicular to the illustration [5]. In the initial stages, rims form at the edges of the film as it retracts, while a groove develops at the GBs where the film eventually breaks. Afterward, every grain evolves following the SSD dynamic of a single crystal, resulting in a number of final islands equal to the number of initial grains. In the central grain at $t=0.25$, the typical material depletion in the center due to surface diffusion [1] can be observed, leading to a dent in the surface. For grains with large aspect ratios, as obtained after recrystallization, impingement of single-crystal patches, or abnormal grain growth, material depletion can cause breakpoints within grain centers, leading to more islands than initial grains.
@@
Figure 3: Simulations (3D) of morphological evolution and SSD in polycrystalline thin films with an idealized initial microstructure. (a) Perspective view of the equilibrium configuration for thin films with varying aspect ratio, except for $r^{\rm 3D}=16$ for which an intermediate stage is reported. (b) Representative stages of SSD for grains with $r^{\rm 3D}=8$ (top view). The timescale is expressed relative to the breakup time. (c) Representative stages of SSD for grains with $r^{\rm 3D}=32$ (top view). The timescale is set as in panel (b).
@@
The well-known Young’s law can be applied to calculate the contact angles at the GB. There, the grooves exhibit an (equilibrium) angle $\theta$ w.r.t the planar surface of the initial film that satisfies the identity $2\gamma_{\rm s}\sin\theta=\gamma_{\rm b}$ where $\gamma_{\rm s}$ and $\gamma_{\rm b}$ are the (isotropic) surface energy density of the crystal surface and the boundary, respectively. With the chosen interfacial energies ($\gamma_{\rm s}=\gamma_{\rm b}=0.5$), it is $\theta=\pi/6$ and is well reproduced by the MPF simulations. For a GB between two infinitely extended grains ($r^{\rm 2D}\rightarrow\infty$), the depth of the grooves monotonously increases over time [8], thus inherently leading to breakup. For a finite aspect ratio, however, surface diffusion and thus grooving at the GBs stops when the equilibrium shape of the surface of the grain(s) is reached [10]. For isotropic surface energy, the equilibrium shape corresponds to a circular profile that satisfies the equilibrium angle $\theta$ at the groove. By simulating a periodic system with uniformly distributed grains and systematically increasing their aspect ratio, we identify a critical aspect ratio $r_{\rm c}^{\rm 2D}$ above which SSD occurs ($r^{\rm 2D}>r_{\rm c}^{\rm 2D}$). It results approximately $r_{\rm c}\approx 9$ for the chosen parameters, as illustrated in Fig. 1c. We remark that this occurs when the (diffuse) interface meets the substrate.
@@
Following Ref. [10], $r_{\rm c}^{\rm 2D}$ can be independently estimated to validate the simulation results and the underlying framework. For an interface of infinitesimal thickness, $r_{\rm c}^{\rm 2D}$ is determined by considering mass conservation, i.e., equating the initial grain area to the area of a circular cap with contact angle $\theta$ and base $W$, i.e. $A_{\rm G}=A_{\rm C}$, with $A_{\rm G}=WH$ and $A_{\rm C}=W^{2}(2\sin\theta)^{-2}(\theta-\sin\theta\cos\theta)$. For an interface with a finite effective thickness, where breakup occurs when the groove reaches a distance $D$ from the substrate, an additional area $A_{D}=WD$ must be accounted for. The mass conservation condition then reads $A_{\rm G}=A_{\rm C}+A_{\rm D}$. Expressing lengths as fractions of the initial film thickness ($W=rH$ and $D=dH$, implying $d<1$) and solving for the (critical) aspect ratio, we get
@@
For $d=0$ and $\theta=\pi/6$ this results into $r_{\rm c}\approx 11$. By setting $d=\epsilon/H=0.2$, we get $r_{\rm c}\approx 8.8$, which is in excellent agreement with the simulation results. This agreement also supports the use of Eq. (5) to predict the sharp interface limit, $d=0$.
We extend here the previously outlined geometrical and energetic arguments to three dimensions. In an idealized scenario where regular $n$-gon shaped grains of uniform size perfectly tile the substrate [10, 30] (see Fig. 2b for $n=6$) the aspect ratio can be defined as $r^{\rm 3D}=W/H$ where $W=2a$ and $a$ is the apothem of the $n$-gon. The volume of each grain with height $H$ is then given by $V_{\rm G}=A_{n}H$ with the base area $A_{n}=n\tan(\pi/n)W^{2}/4$. As an ideal equilibrium condition, the system is assumed to evolve toward a spherical cap truncated by the $n$-gon (see Fig. 2c). As derived in the Supplementary Material (SM), the corresponding integral expression for its volume is
@@
$\tilde{\rho}=\rho/a$, $\tilde{R}=R/a=1/\cos(\pi/n)$, $\vartheta_{n}(\tilde{\rho})=\pi/n$ if $\tilde{\rho}<1$ and $\pi/n-\arccos(\tilde{\rho}^{-1})$ otherwise. By imposing mass conservation, $V_{\rm G}=V_{\rm C}+V_{\rm D}$ with $V_{\rm D}=A_{n}D$ and $D=dH$ in analogy to Eq. (5), the critical aspect ratio results
@@
Fig. 2d illustrates the dependence of the critical aspect ratios, Eqs. (5) and (8) for regular hexagonal grains ($n=6$), on $\theta$. Note that $r_{\rm c}^{\rm 3D}$ increases with $n$, indicating that grains with fewer edges undergo SSD earlier than those with more edges. This suggests that sharp corners promote film rupture. Note that from the equations above, $\tilde{z}=0$ is ensured at the junction where $\tilde{\rho}=\tilde{R}$. Since the (spherical) elevation decreases with radius, an enhanced grooving at junctions as observed in previous works [34, 30] is naturally accounted for. The elevation of the junction (J) and of the GB at the midpoint between junctions can be expressed as (see SM)
$\begin{split}d_{\rm J}&=H-z_{n}(\theta,R),\\ d_{\rm GB}&=\frac{W}{2}\left(\frac{1}{\tan{\theta_{c}}}-\sqrt{\frac{1}{\sin^{2}{\theta_{c}}}-\frac{1}{\cos^{2}({\pi}/{n})}}\right)+d_{\rm J}.\end{split}$ (9)
Figure 4: SSD of square polycrystalline patches with an arbitrary internal grain structure. The two panels (a) and (b) show representative stages of the SSD evolution with two different realizations of the initial conditions featuring different spreads in grain sizes and shapes but analogous qualitative behaviors. The timescales are arbitrarily expressed relative to the breakup time (similar in both the illustrated cases). Note that panel (b) illustrates the evolution until the very late stages.
MPF simulations of SSD of the idealized initial configurations as in Figs. 2b are presented in Fig. 3a varying $r^{\rm 3D}$. For relatively small aspect ratios, the morphology of grains at equilibrium is well described by spherical caps truncated by hexagons ($n=6$ in Fig. 2c). $d_{\rm J}$ and $d_{\rm GB}$ decrease with increasing $r^{\rm 3D}$ and correspond well to the prediction of Eq. (9). For instance, considering the case of $r^{\rm 3D}=5.5$, the predicted elevations normalised to the initial film thickness are $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.64$ and the simulated normalised elevations measured at the nominal interface between vapor phase and grains ($\phi_{1}=0.5$) are $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.65$. Moreover, a critical aspect ratio of $r^{\rm 3D}_{\rm c}\approx 6.7$ is obtained from simulations (c.f. Fig. 3a). Eq. (8), for the considered scenario where $\theta=\pi/6$, delivers an estimation of $r_{\rm c}^{\rm 3D}\approx 8.7$ for $d=0$ and $r_{\rm c}^{\rm 3D}\approx 6.8$ considering the diffuse interface of MPF 3D simulations with $d=\epsilon/H=0.22$ (see also Fig. 2d), in excellent agreement with the result of simulations. On the one hand, we can conclude that the MPF framework well reproduces estimation based on interface/surface energy minimization. On the other hand, predictions of Eqs. (8) and (9) are confirmed by (independent) numerical simulations. They can be readily exploited to provide initial estimates of the onset of SSD in polycrystalline thin films, given the aspect ratio of the grains. We note, however, that localized inhomogeneities and defects typically occur in polycrystals, which can trigger SSD but with a highly heterogeneous character.

### v2

@@ Abstract
Solid-state dewetting is the process by which thin solid films break up and retract on a substrate, forming nanostructures. While dewetting of single-crystalline films is understood as a surface-energy-driven process mediated by surface diffusion, polycrystalline films exhibit additional complexity due to the presence of grain boundaries. Most theoretical and computational studies have focused on single-crystalline dewetting. Here, we present the application of the grandpotential multi-phase-field model to the dewetting of thin polycrystalline films in three dimensions, reproducing the key phenomenology of this process. By considering isotropic interface/surface energy, we illustrate its consistency with predictions based on energetic arguments and the morphological evolution towards equilibrium. We also provide novel analytical criteria for the onset of three-dimensional dewetting, serving as fundamental theoretical benchmarks, and highlight the critical role of triple junctions. Moreover, we unveil the dewetting behavior of polycrystalline patches, extending the scenarios of their single-crystalline counterparts.
Thin films, grain boundaries, surface diffusion, phase-field model
@@
Thin crystalline films are metastable due to their high surface-to-volume ratio. At high temperatures below melting, they typically break up and form three-dimensional agglomerates in a process known as solid-state dewetting (SSD) [1]. While detrimental to planar architectures, this phenomenon offers a unique kinetic pathway for the self-assembly of complex nanostructures, as clearly demonstrated for single-crystalline thin films. [2, 3, 4, 5]. In these systems, the dynamics are predominantly governed by surface diffusion [1].
@@
Several technologically relevant thin films are polycrystalline [6, 7]: they consist of crystalline domains (grains) with varying crystallographic orientations, separated by interfaces (grain boundaries, GBs). Their morphological evolution is influenced not only by surface diffusion, as in single crystals, but also by additional effects associated with GBs [8, 9, 10, 11]. Arguably, the most relevant one is that surface grooves develop at GBs. Moreover, GBs migrate through a mechanism resembling attachment-detachment kinetics, with a velocity governed largely by capillarity (mean curvature flow) and potentially influenced by elasticity effects [12, 13].
@@
Modeling and simulations have been pivotal in understanding and tailoring single-crystalline SSD [14, 4, 5]. Studies on GB grooving and few-grain microstructures provided valuable insights [15, 16, 17, 18, 19], yet they remain limited in capturing the coupled GB migration and surface diffusion dynamics over many grains and long timescales. Phase-field (PF) models [20, 21, 22, 23] have emerged as a leading framework for SSD studies. They realize an implicit description of geometries via smooth order parameter(s), which naturally handles topological changes, complex geometries, and three-dimensionality, all key aspects of SSD [24, 4, 5, 25, 26]. Convenient PF models coupling diffusion, phase transformation, and interface migration have been proposed [27, 28, 18, 29, 30]. Recent work has examined interesting aspects of polycrystalline thin film evolution, including the impact of thermal grooving on grain growth and its subsequent stagnation [31, 32]. First studies of polycrystalline dewetting have also been proposed [33], although with an approximated treatment of evolving morphologies and surface diffusion. An up and coming PF formulation, namely the grand-potential multi PF model [27, 28, 34] (hereafter MPF), offers a general descriptions capturing the dynamics of both conserved and non-conserved order parameters, thus ideal for polycrystalline SSD scenarios which include (surface and bulk) diffusion as well as interface (GB) migration. Although this scenario has not been explored in detail, a recent implementation promisingly showed key phenomenology concerning thermal grooving at one GB [29].
We present here the application of this MPF model to studying SSD in polycrystalline thin films. After outlining the model’s key features, we demonstrate its capabilities. We examine SSD onset through numerical simulations and benchmark its outcome against newly derived equations that predict critical grain aspect ratios and related features. Finally, we apply the MPF model to prototype polycrystalline films and to the relevant case of patches, highlighting similarities and differences with single-crystalline counterparts.
@@
We employ the approach introduced in [29], adapted here to systems dominated by surface/interface diffusion and interface migration, while neglecting bulk diffusion. We also consider grains of the same pure material and neglect any elastic effects arising from GB migration. Order parameters $\{\phi_{1}(\mathbf{r}),\dots,\phi_{N}(\mathbf{r})\}\equiv$$\boldsymbol{\varphi}$ akin to smooth characteristic functions track different domains/phases. Global concentration fields for $K$ elements $c_{i}(\mathbf{r})$ account for material transport. They can be related to the composition of each phase ($c_{i\alpha}$) via interpolation $c_{i}(\mathbf{r})=\sum_{\alpha=1}^{N}\phi_{\alpha}(\mathbf{r})c_{i\alpha}$ (spatial dependence will be omitted for brevity). To describe the chemical interactions within the MPF system, we define a set of K1K-1 independent chemical potential fields, $\{\mu_{1},\dots,\mu_{K-1}\}\equiv\boldsymbol{\mu}$, corresponding to the concentration fields, which distinguish between the solid film, the substrate, and the surrounding vacuum. The grand potential functional reads
[…]
@@
$f_{\rm I}=\sum_{\alpha}^{N}\sum_{\beta=\alpha+1}^{N}\gamma_{\alpha\beta}\left(\epsilon\nabla\phi_{\alpha}\cdot\nabla\phi_{\beta}+\frac{16\mathcal{G}(\boldsymbol{\varphi})\phi_{\alpha}\phi_{\beta}}{\pi^{2}\epsilon}\right).$ (2)
$f_{\rm I}$ encodes the interfacial energy with $\epsilon$ the interface thickness, $\mathcal{G}(\boldsymbol{\varphi})=1$ if $\sum_{\alpha}\phi_{\alpha}=1$ and $\phi_{\alpha}\geq 0,\ \forall\alpha$, while $\mathcal{G}(\boldsymbol{\varphi})=\infty$ otherwise. The function $\psi(\boldsymbol{\varphi},\boldsymbol{\mu})$ corresponds to the grand-potential density
@@
We consider here a parabolic free energy density $f_{i\alpha}=A_{i}^{\rm chem}(c_{i\alpha}-c_{i\alpha}^{\text{eq}})^{2}$ depending on the equilibrium concentrations $c_{i\alpha}^{\text{eq}}$ and constants $A_{i}^{\rm chem}$.
@@
with $m_{\alpha\beta}$ and $M_{\alpha\beta}$ the mobility and diffusivity at the $\alpha\beta$-interface, respectively. The timescale will be expressed in arbitrary units. Sharp interface limits can be found in [29, 35]. Simulations are obtained by numerically solving Eqs. (4) using a finite difference integration scheme, implemented in the modular software package PACE3D [36]. Periodic boundary conditions are applied in the in-plane directions, while no-flux Neumann conditions are imposed at the top and bottom boundaries.
@@
Figure 1: Modeling and 2D simulation of polycrystalline SSD. (a) Scheme of the considered geometries and labeling of the phases as considered in the MPF model. The grain aspect ratio is defined as $r^{\rm 2D}=W/H$. (b) SSD simulations ($\epsilon=0.2H$) of a thin film with three grains. The central one has an aspect ratio of $r^{\rm 2D}=25$. The two at the border of the film have both $r^{\rm 2D}=12.5$. The timescale is relative to the one needed to reach equilibrium. The inset shows the vapour-solid-substrate trijunction. (c) MPF simulations of SSD with varying $r^{\rm 2D}$ for a film ideally formed by grains of the same size. A critical aspect ratio of $r^{\rm 2D}_{\rm c}\approx 9$ is obtained in simulations.
@@
Figure 2: Geometrical parametrization and estimation of the critical aspect ratios. (a) Schematics of 2D equilibrium morphology. The dashed line illustrates the thickness ($H$) of the conformal initial film. (b) Schematics (top view) of the idealized setting of hexagonal grains in 3D with equal size. (c) Spherical caps truncated by $n$-gons. (d) $r_{\rm c}(\theta)$ in 2D and 3D as predicted by Eqs. (5) and (8) for $d=0$ (sharp interface limit) and $d>0$ as in simulations. Symbols show $r_{\rm c}$ obtained by MPF simulations with $d=0.22$ (2D) and $d=0.25$ (3D). The inset illustrates the dependence of $r_{\rm c}$ on $d$ for $\theta=\pi/6$ (2D), as well as the average standard deviation of simulation results ($\pm 0.25$).
@@
In our polycrystalline thin film model, the order parameters $\phi_{\alpha}$ track the substrate, the vapor phases, and the different grains within the film. Fig. 1a shows an illustration of a polycrystalline film composed of three grains, where we label the vacuum phase $\alpha=1$, the substrate $\alpha=N$, and the grains $1<\alpha<N$. We refer to the ratio of width to height $r^{\rm 2D}=W/H$ as the aspect ratio of the 2D grains. We use a two-concentration-field model for the substrate and the film with $c_{\text{substrate},N}^{\text{eq}}=1$ and $c_{\text{grain},\alpha}^{\text{eq}}=1$. We set $A_{\text{substrate}}^{\rm chem}=50$ and $A_{\text{grain}}^{\rm chem}=5$ [35]. Interfacial (surface) diffusion is only allowed at the film-vapor interface via $M_{\alpha\beta}=\delta_{1\beta}$. We set $\gamma_{\alpha\beta}=0.5$ for all interfaces except otherwise specified. GB mobilities are set to $10^{-3}$ while all other mobilities are set to $1$. This choice allows for inspecting dewetting regimes with relatively slow interface motion. It does not prevent grain growth, as further illustrated below. Parameters match those used in Ref. [37] for modeling (single-crystalline) nickel coarsening in solid oxide fuel cells. We indeed neglect interface-energy and mobility anisotropy, which are out of scope for the present discussion.
@@
The simulation of SSD for a 2D three-grains film is shown in Fig. 1b. In the initial stages, rims form at the edges of the film as it retracts, while a groove develops at the GBs where the film eventually breaks. Afterward, each grain evolves according to the SSD dynamics of a single crystal. In the central grain at $t=0.25$, the typical material depletion in the center due to surface diffusion [1] can be observed, leading to a dent in the surface. For grains with large aspect ratios, as obtained after recrystallization, impingement of single-crystal patches, or abnormal grain growth, material depletion can cause breakpoints within grain centers, leading to more islands than initial grains.
@@
Figure 3: Simulations (3D) of morphological evolution and SSD in polycrystalline thin films with an idealized initial microstructure ($\epsilon=0.22H$). (a) Perspective view of the equilibrium configuration for thin films with varying aspect ratio, except for $r^{\rm 3D}=16$ for which an intermediate stage is reported. (b) Representative stages of SSD for grains with $r^{\rm 3D}=8$ (top view). The timescale is expressed relative to the breakup time. (c) Representative stages of SSD for grains with $r^{\rm 3D}=32$ (top view). The timescale is set as in panel (b).
@@
The grooves exhibit an (equilibrium) angle $\theta$ w.r.t the planar surface of the initial film that satisfies the identity $2\gamma_{\rm s}\sin\theta=\gamma_{\rm b}$ where $\gamma_{\rm s}$ and $\gamma_{\rm b}$ are the (isotropic) surface energy density of the crystal surface and the boundary, respectively. In this simulation $\gamma_{\rm s}=\gamma_{\rm b}=0.5$, i.e. $\theta=\pi/6$, well reproduced by the MPF simulations. For a GB between two infinitely extended grains ($r^{\rm 2D}\rightarrow\infty$), the depth of the grooves monotonously increases over time [8], thus inherently leading to breakup. For a finite aspect ratio, however, surface diffusion and thus grooving at the GBs stops when the equilibrium shape of the surface of the grain(s) is reached [10]. For isotropic surface energy, the equilibrium shape corresponds to a circular profile that satisfies the equilibrium angle $\theta$ at the groove. By simulating a periodic system with uniformly distributed grains and systematically increasing their aspect ratio, we identify a critical aspect ratio $r_{\rm c}^{\rm 2D}$ above which SSD occurs ($r^{\rm 2D}>r_{\rm c}^{\rm 2D}$). It results approximately $r_{\rm c}\approx 9$ for the chosen parameters, as illustrated in Fig. 1c. We remark that this occurs when the (diffuse) interface meets the substrate.
@@
Values of $r_{\rm c}^{\rm 2D}$ can be independently estimated by considering mass conservation [10], i.e., equating the initial grain area to the area of a circular cap with contact angle $\theta$ and base $W$, i.e. $A_{\rm G}=A_{\rm C}$, with $A_{\rm G}=WH$ and $A_{\rm C}=W^{2}(2\sin\theta)^{-2}(\theta-\sin\theta\cos\theta)$. We also introduce an additional area $A_{D}=WD$ to account for breakup occurring when the groove reaches a distance $D$ from the substrate, namely in the presence of interfaces/surfaces with a finite effective thickness. The mass conservation condition thus reads $A_{\rm G}=A_{\rm C}+A_{\rm D}$. Expressing lengths as fractions of the initial film thickness ($W=rH$ and $D=dH$, implying $d<1$) and solving for the (critical) aspect ratio, we get
@@
For $d=0$ and $\theta=\pi/6$ this results into $r_{\rm c}\approx 11$. By setting $d=\epsilon/H=0.2$, we get $r_{\rm c}\approx 8.8$, which is in excellent agreement with the simulation results. This holds upon varying $\theta$, which can be controlled by adjusting the solid–vapor interfacial energy ($\gamma_{\rm s}$, represented by $\gamma_{\alpha 1}$ in the MPF model), and $d$, as illustrated in Fig. 2d.
For an idealized 3D case of uniformly sized, regular $n$-gonal grains perfectly tiling the substrate [10, 30] (Fig. 2b, $n=6$), the aspect ratio is defined as $r^{\rm 3D}=W/H$ where $W=2a$ and $a$ is the apothem of the $n$-gon. The volume of each grain is given by $V_{\rm G}=A_{n}H$ with the base area $A_{n}=n\tan(\pi/n)W^{2}/4$. The system is assumed to evolve toward a spherical cap truncated by the $n$-gon (see Fig. 2c). As derived in the Supplementary Material (SM), the corresponding integral expression for its volume is
@@
$\tilde{\rho}=\rho/a$, $\tilde{R}=R/a=1/\cos(\pi/n)$, $\vartheta_{n}(\tilde{\rho})=\pi/n$ if $\tilde{\rho}<1$ and $\pi/n-\arccos(\tilde{\rho}^{-1})$ otherwise. By imposing mass conservation, $V_{\rm G}=V_{\rm C}+V_{\rm D}$ with $V_{\rm D}=A_{n}D$ and $D=dH$ in analogy to Eq. (5), we then find the critical aspect ratio
@@
It increases with $n$, indicating that grains with fewer edges undergo SSD earlier, suggesting that sharp corners promote film rupture. From the equations above, $\tilde{z}=0$ at the junction where $\tilde{\rho}=\tilde{R}$; since the (spherical) elevation decreases with radius, enhanced grooving at junctions, observed in previous works [38, 30], is naturally captured. The elevation of the junction (J) and of the GB at the midpoint between junctions can be expressed as (see SM)
[…]
Figure 4: SSD of square polycrystalline patches with an arbitrary internal grain structure. The two panels (a) and (b) show representative stages of the SSD with two different initial conditions obtained through a Voronoi construction over randomly generated points. The resulting domains are then randomly assigned to different phase fields. These initial conditions show significantly different variability of the numbers of grains per side, which are $\sim 7.25\pm 0.69$ in (a) and $\sim 4.75\pm 3.69$ (b). Timescale is given relative to the polycrystalline patch breakup time (similar in both cases). Panel (b) shows the evolution up to the late stages. (c) SSD of a single-crystalline patch with the same size and aspect ratio as in (a) and (b).
MPF three-dimensional simulations of SSD of the idealized initial configurations in Figs. 2b are presented in Fig. 3a, varying $r^{\rm 3D}$. For relatively small aspect ratios, grain morphologies at equilibrium are well described by hexagonally truncated spherical caps ($n=6$, Fig. 2c). Both $d_{\rm J}$ and $d_{\rm GB}$ decrease with increasing $r^{\rm 3D}$ and agree with Eq. (9). For instance, at $r^{\rm 3D}=5.5$, simulations give $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.65$, in almost perfect agreement with predictions of Eq. (9), namely $d_{\rm J}/H\approx 0.37$ and $d_{\rm GB}/H\approx 0.64$. The proposed Eq. (8) further estimates $r_{\rm c}^{\rm 3D}\approx 8.7$ for $d=0$ and $r_{\rm c}^{\rm 3D}\approx 6.8$ for $d=\epsilon/H=0.22$, well matching the simulated critical aspect ratio $r_{\rm c}^{\rm 3D}\approx 6.7$ (Fig. 3a). This agreement holds for other $\theta$ values too, as illustrated in Fig. 2d. The MPF framework is thus benchmarked against theoretical predictions from Eqs. (8) and (9), which also independently enable the estimation of the SSD onset in polycrystalline thin films from the grain aspect ratio. We note, however, that localized inhomogeneities and defects may trigger SSD with pronounced heterogeneity.